OpenStudy (anonymous):

solve... 4x(x-6)-5x(x-5)=-2

OpenStudy (anonymous):

4x^2 -24x -5x^2 + 25x +2 = 0 x^2 -x -2 = 0 (x-2)(x+1) = 0 x =2 or x=-1

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ok how about a medal?